OpenStudy (anonymous):

what is b to the 5 power for b = -2? show work!!

OpenStudy (anonymous):

(-2)^ 5 = -2*-2*-2*-2*-2

OpenStudy (amistre64):

exponential notation is just shorthand for mulitplying alot of times by the same number

OpenStudy (amistre64):

b^5 = b*b*b*b*b

OpenStudy (amistre64):

multiplication itself is shorhand notation for adding a value to its self alot of times :) 4b = b+b+b+b

OpenStudy (amistre64):

and numbers themselves are shorthand notation for adding up alot of 1s 3 = 1+1+1 :)

OpenStudy (radar):

In the case of a negative number (as above) if the exponent is odd the result will be negative, if it is even the result will be positive.

OpenStudy (anonymous):

give ma a answer and work it out!!!

OpenStudy (amistre64):

jimmy worked it out; you are adept enough to get the specific answer from that i believe

OpenStudy (anonymous):

i dont understand that so can you work it out and give me the answer!!

OpenStudy (amistre64):

i can, but i dont believe it would benefit you in any way ... srry

OpenStudy (anonymous):

"exponential notation is just shorthand for mulitplying alot of times by the same number" Well this is true until you raise a number to a non-integer or negative power.

OpenStudy (amistre64):

negative powers relate to division; which is just multiplication in disguise :)

OpenStudy (anonymous):

Yes but your statement is no longer correct. You would be multiplying by the inverse of the number not by the number.

OpenStudy (amistre64):

good notation is the bridge to understanding i believe ;)

OpenStudy (anonymous):

To better understand "raising a number to a power" you must learn the properties of the exponential function.